# 1pt plus 3pt minus 2pt

### v1

@@ IV-D Progress is used to signify the completion rate of a task
The results of online study two are shown in Figure 4. Each line represents the average progress or scalar feedback of 18 participants under a certain scenario. While scalar feedback and progress can both describe the quality of a demonstration, progress and scalar feedback focus on different information.
Progress describes the degree to which a task has been completed. Progress was low at the beginning, increased as the task was completed more, stayed the same or became lower while the robot was not acting towards finishing the task (see Failure before spinning, Unaware after missing the stirrer, and Corrected before re-picking the stirrer). A higher value of progress indicates more successful task completion. For instance, in the Unaware scenario, the progress at the final step is greater than in the Failure scenario but lower than in the Perfect scenario, due to 5 out of 6 sub-tasks being completed in the Unaware, as opposed to 1 out of 6 in the Failure. We calculate Pearson corrections between the progress function from participants and an accumulative progress function created by true labels (+1 for completing part of the task, 0 for not completing anything, and -1 for backtracking) for each scenario. The correlations are 0.983 for Perfect scenario ($p<0.001$), 0.982 for Imperfect scenario ($p<0.001$), 0.974 for Unaware scenario ($p<0.001$), 0.977 for Corrected scenario ($p<0.001$), and 0.895 for Failure scenario ($p<0.001$). H2 is supported.
@@ IV-E Progress indicates if a task is complete
We showed average progress and average scalar feedback at the last step for all participants and statistic analysis results between each scenario and Perfect scenario in Table I. We found that progress could correctly indicate if the is complete. The task has been completed in three scenarios, the Perfect, Imperfect, and Corrected. The average progress of three completed scenarios at the last step is all about 90, and there was no evidence showing that there is any difference between the other two and the Perfect case ($p=0.71,BF=0.34$ for Imperfect, and $p=0.62,BF=0.33$). For the Unaware and Failure scenarios, there is strong evidence indicating that the task was not completed ($avg=61.3,p<0.01,BF=98.75$ for Unaware, and $avg=23.4,p<0.01,BF>1000$ for Failure). H3. is supported. This can not be captured by scalar feedback.
IV-F Progress is robust and more consistent to sub-optimal demonstrations
In the Imperfect scenario, when the robot dropped the cup onto the table, the progress remained at a similar level and did not affect the eventual progress. The conclusion holds the same in the Corrected scenario. The progress only changed slightly when the robot missed the stirrer and started spinning unproductively, and progress at the end is similar to Perfect. H4 is supported. Scalar feedback, on the other hand, changed dramatically in all these cases. Participants used scalar feedback to indicate if a demonstration was clean and good, but the possibility that the errors were ”harmless explorations” or ”fixable mistakes” is not captured by scalar feedback.
Moreover, for every scenario other than Perfect, progress has a lower average standard deviation in each scenario compared to scalar feedback. The average standard deviation for Imperfect: 16.1 for progress and 22.5 for scalar feedback ($p=0.02$, $BF=3.19$), Unaware: 16.8 for progress and 21.1 for scalar feedback ($p=0.06$, $BF=1.41$), Corrected 16.1 for progress and 20.3 for scalar feedback ($p=0.06$, $BF=1.39$), and Failure: 15.6 for progress and 28.3 for scalar feedback ($p<0.001$, $BF>1000$). H5 is supported.
@@ V Progress with Self-Provided Demonstrations from Non-experts
While progress can be effective in being used alone with expert demonstrations, using progress to annotate non-expert demonstrations may still be challenging. We expected most hypotheses will hold and progress will be more consistent across participants than scalar feedback if demonstrations are imperfect and multi-modal We conducted a public space study to validate the applicability of progress with self-provided demonstrations from non-experts. We recruited 40 participants to provide demonstrations, progress, and scalar feedback in an ice cream topping-adding task. Participants were recruited from the atrium of a university building and the overall participation time was about 15 minutes. This work is approved by the Institutional Review Board and all data collected was anonymous. We released all data we collected from our public space study as a dataset, along with the example scripts that read the data from files. The dataset is available at: https://github.com/TeachingwithProgress/Non-Expert_Demonstrations
@@ V-A Experiment Setup
The study was settled in the lobby of a university building. Each participant was asked to first give one demonstration and then watch a replay of the demonstration. The task we asked participants to demonstrate is an ice cream topping-adding task. The goal for participants is to pick up a topping from a shelf and pour the topping into an ice cream via teleoperating a robot arm. The shelf is located on the left side of the workspace, and there are four toppings available which are located at four locations. The participants controlled the arm by using an Xbox controller. The arm was a Kinova Gen 3 Lite arm with six DoF. The setup and the workspace are shown in Figure 1. We recorded the demonstrations in 5 HZ. During the replay, the arm would stop after finishing replaying 10% of the frames, and we would ask participants for one progress and one scalar feedback for the replayed partial trajectory.
@@ V-B Experiment Procedure
We recruited participants by asking people who walked by our setup. Of the 40 participants, 22 participants were male, 14 participants were female, and 4 participants preferred not to say. If the participants agreed to join the study, we then asked them if they were familiar with teleoperating a robot using a controller to ensure that they were not experts, and asked them to fill out a consent form to ensure they were eligible for the study. Then we introduced them to the ice cream topping adding task and how to use an Xbox controller to teleoperate the arm. Participants could have up to 3 minutes to practice the task. After practicing, participants were asked to give demonstrations. Each participant only had one chance and could not retry the task even if they failed in the middle due to imperfectly demonstrating. Once participants had finished giving a demonstration, they would inform the experimenter that the demonstration was finished. Then, the experimenter would stop recording and introduce participants to the replay evaluation task. Before starting replaying, the experimenter would introduce progress and scalar feedback in detail to reduce the difference in understanding of the signals among participants. During the replay, the arm would stop ten times. For each time, the experimenter would briefly explain progress and scalar feedback again, and ask participants for two signals in a random order.
V-C Non-Expert Demonstrations Are Multi-Policy and Noisy but Mostly Succeed
@@
We collected 40 demonstrations from 40 participants. All trajectories are visualized in Figure 5. We also visualized the locations of the objects used in the experiment. The blue trajectories are successful demonstrations and the orange trajectories are failed demonstrations. We found that non-expert demonstrations are noisy and contain a variety of policies, even though 34 out of 40 are ultimately successful. This suggests that policies can be both successful and sub-optimal, which supports our intuition: assessing the quality of demonstrations by comparing is likely to lose information if there are many ”good enough” policies. We also observed that the noise in the demonstrations is not only teleoperating errors but also explorations. For instance, we observed that many participants slightly shook the topping jar to test if the gripper firmly held the jar when picking the jar, or poured a few toppings out first to see if the jar was right above the ice cream when pouring the toppings. This highlights the importance of detecting unproductive behaviors, and suggests that noisy demonstrations from partially trained agents or perfect demonstrations with injected noise are inappropriate approximations of noisy human demonstrations: human demonstrations are “noisy” in a meaningful way. For the six failed demonstrations, [ESS: [x]] of them were because the topping jar was accidentally dropped while reaching the ice cream, [ESS: [x]] of them were because the gripper did not successfully pick the topping jar. The most common faulty cases are similar to the Failure scenario and the Unaware scenario we used in our online study, which confirmed the validity of the design of our online study.
@@ V-D Progress Indicates Task Completions and is more consistent across participants than scalar feedback
We collected progress and scalar feedback from 34 participants, of these, two participants’ progress and scalar feedback were excluded since all progress and scalar feedback they provided were 100, indicating that they misunderstood the task or were deliberately giving incorrect feedback. The results are shown in Figure 6. We plot the average progress and scalar feedback for all successful demonstrations, and progress and scalar individually for all failure demonstrations. We are not able to determine if progress from participants correctly describes the task completion rate since we do not have ground truth labels, but average progress for successful demonstrations did start from low and increased as the demonstrations were reproduced which is a strong indication that progress correlates with task completion rates.
We successfully determined if a demonstration is a successful demonstration by only looking at progress at the last step. We use 90 as the divide value, which is the average progress at the last step for successful demonstrations in our online study. If progress at the last step is less than 90, the task is incomplete, otherwise the task is completed. For all failed demonstrations, the end progress is less than 90 and the average progress at the end for successful demonstrations is 98.8. H3. is supported. Moreover, progress is significantly more consistent across participants even though their demonstrations are multiple policies and in different quality ($avg\_std_{progress}=11.6$, $avg\_std_{scalarfeedback}=25.5$, $p<0.001$, $BF>1000$). H5. is supported.
@@ V-E Progress allows the awareness of backtracking
Figure 7: 3D trajectories of one participant’s demonstration. The trajectories are colored using scalar feedback or progress. The brown circle represents the start point. The green start represents the endpoint. The blue square is the position of the topping jar. The participant reset the arm to the start position after demonstrating a few steps to adjust the grip point.
As mentioned in subsection V-C, participants would do explorations while demonstrating difficult parts of the task. We also observed that some participants went back a few steps to adjust the grip pose or the pouring position. For instance, one participant failed to pour the toppings into the ice cream because the grasp point was not optimal. Then the participant decided to reset the arm to the initial position, demonstrated the entire task from the beginning again, and succeeded. We plot the trajectory for that demonstration along with progress and scalar feedback, shown in Figure 7. As indicated by progress, the arm was reset to the initial position at step 4 and the task was successfully completed afterwards.
@@ VI Discussion
In this work, we proposed a novel teaching signal, progress, and conducted three different studies to show the usefulness of progress across a variety of scenarios. We showed that giving progress is not hard and progress carries information that is beyond scalar feedback and other teaching signals in prior work, such as indicating task completion and harmless explorations. We expect that progress will be effective in many applications other than just using alone with demonstrations.
Prevent Reward Hacking Progress could be a powerful signal to prevent reward hacking. Reward hacking is a phenomenon that a learning agent learns to achieve high rewards by performing unintended actions instead of finishing the task. For instance, a cleaning robot gets a +1 reward every time when it cleans a room. The robot, instead of cleaning one room and going to the next room, repetitively ejects dirt in a room then cleans it, and thus achieves high rewards. Using progress, we can easily identify that the robot is not advancing towards task completion.
Inverse Reinforcement Learning We observed a similarity between progress functions and reward functions and expect that a progress function might be a representation of a reward function. We trained a reward function using 40 demonstrations we collected using AIRL , and calculated the rewards for each demonstration. We then calculated the Pearson correlation between the rewards of the demonstration and the progress, as well as the correlation between the rewards and the scalar feedback from each participant in Figure 8. We found that progress is strongly correlated with the learned reward function ($avg_{r}=0.83$) and the average correlation is significantly higher ($p<0.001$, $BF>1000$ )) than scalar feedback ($avg_{r}=0.19$).
Data Filtering and Ranking Progress could also be used to rank demonstrations. For example, a demonstration with a progress of 60 at the end should be ranked lower than a demonstration with a progress of 90. Moreover, progress can be applied as a data filter especially when the demonstrations are sub-optimal. For example, if a participant made a mistake in the middle of a demonstration and then fixed it, we could filter out the part with negative delta progress labels, and keep the rest since we know the task is completed. With scalar feedback, we would only know there was a mistake by using scalar feedback.
A key area for future work is to compare the model performance between the model trained using progress and the model trained using scalar feedback. Training a reliable model from limited non-expert demonstrations and non-expert annotations is challenging but future work could expand our data with more demonstrations along with more types of human teaching signals such as preference and ranking.
@@ VII Conclusion
In conclusion, our proposed teaching signal progress could be used to describe completion degrees of a task, indicate if a task is complete, and detect unproductive but harmless behaviors, while giving progress does not require extra workload and time compared to giving scalar feedback. We collected 40 non-expert demonstrations along with progress and scalar feedback, and released them as a dataset. We found that non-expert demonstrations are multi-policy and mostly successful, while noisy in a meaningful way. Our work suggests that progress is information-rich, and is worth more attention to develop new methods to effectively leverage the novel information from progress.

### v2

@@ IV-D Progress is used to signify the completion rate of a task
The results of online study II are shown in Figure 4. Each line represents the average progress or scalar feedback of 18 participants under a certain scenario. While scalar feedback and progress can both describe the quality of a demonstration, progress and scalar feedback focus on different information.
Progress describes the degree to what extent a task has been completed. Progress was low at the beginning, increased as the task was completed more, stayed the same or became lower while the robot was not acting towards finishing the task (see Failure before spinning, Unaware after missing the stirrer, and Corrected before re-picking the stirrer). A higher value of progress indicates more successful task completion. For instance, in the Unaware scenario, the progress at the final step is greater than in the Failure scenario but lower than in the Perfect scenario, due to 5 out of 6 sub-tasks being completed in the Unaware, as opposed to 1 out of 6 in the Failure. We calculate Pearson corrections between the progress function from participants and an accumulative progress function created by true labels (+1 for completing part of the task, 0 for not completing anything, and -1 for backtracking) for each scenario. The correlations are 0.983 for Perfect scenario ($p<0.001$), 0.982 for Imperfect scenario ($p<0.001$), 0.974 for Unaware scenario ($p<0.001$), 0.977 for Corrected scenario ($p<0.001$), and 0.895 for Failure scenario ($p<0.001$). H2 is supported.
@@ IV-E Progress indicates if a task is complete
We showed average progress and average scalar feedback at the last step for all participants and statistical analysis results between each scenario and the Perfect scenario in Table I. We found that progress could correctly indicate if the task is complete, while the indication of task completion was not captured by scalar feedback. The task has been completed in three scenarios, Perfect, Imperfect, and Corrected. The average progress of three completed scenarios at the last step is about 90, and there was no evidence showing that there is any difference between the other two and the Perfect case ($p=0.714,BF=0.34$ for Imperfect, and $p=0.620,BF=0.33$ for Corrected). For the Unaware and Failure scenarios, there is strong evidence indicating that the task was not completed ($avg=61.3,p<0.001,BF=98.75$ for Unaware, and $avg=23.4,p<0.001,BF>1000$ for Failure). H3 is supported.
IV-F Progress is robust and more consistent to sub-optimality
In the Imperfect scenario, when the robot dropped the cup onto the table, the progress remained at a similar level and did not affect the eventual progress. The conclusion holds the same in the Corrected scenario. The progress only changed slightly when the robot missed the stirrer and started spinning unproductively, and progress at the end is similar to Perfect. Scalar feedback, on the other hand, changed dramatically in all these cases. Participants used scalar feedback to indicate if a demonstration was clean and good, but the possibility that the errors were ”harmless explorations” or ”fixable mistakes” is not captured by scalar feedback. Moreover, for every scenario other than Perfect, progress has a lower average standard deviation in each scenario compared to scalar feedback. The average standard deviation for Imperfect: 16.1 for progress and 22.5 for scalar feedback ($p=0.018$, $BF=3.19$), Unaware: 16.8 for progress and 21.1 for scalar feedback ($p=0.060$, $BF=1.41$), Corrected 16.1 for progress and 20.3 for scalar feedback ($p=0.058$, $BF=1.39$), and Failure: 15.6 for progress and 28.3 for scalar feedback ($p<0.001$, $BF>1000$). H4 is supported.
@@ V Progress with Self-Provided Demonstrations from Non-experts
While progress can be effective in being used along with expert demonstrations, using progress to annotate non-expert demonstrations may still be challenging. We expected most hypotheses will hold. We conducted a public space study to validate the applicability of progress with self-provided demonstrations from non-experts. We recruited 40 participants to provide demonstrations, progress, and scalar feedback in an ice cream topping-adding task. Participants were recruited from the atrium of a university building and the overall participation time was about 15 minutes. This work is approved by the Institutional Review Board and all data collected was anonymous. We released all data we collected from our public space study as a dataset, along with the example scripts that read the data from files. The dataset is available at: https://github.com/TeachingwithProgress/Non-Expert_Demonstrations
@@ V-A Experiment Setup
The study was settled in the lobby of a university building. Each participant was asked to first give one demonstration and then watch a replay of the demonstration. The task we asked participants to demonstrate is an ice cream topping-adding task. The goal for participants is to pick up a topping from a shelf and pour the topping into an ice cream via teleoperating a robot arm. The shelf is located on the right side of the workspace, and there are four toppings available which are located at four locations. The participants controlled the arm by using an Xbox controller. The arm was a Kinova Gen 3 Lite arm with six DoF. The setup and the workspace are shown in Figure 1. We recorded the demonstrations in 5 HZ. During the replay, the arm would stop every 10% of the frames, and we would ask participants for one progress and one scalar feedback for the replayed partial trajectory.
@@ V-B Experiment Procedure
We recruited participants by asking people who walked by our setup. Of the 40 participants, 22 participants were male, 14 participants were female, and 4 participants preferred not to say. If the participants agreed to join the study, we first asked them if they were familiar with robots, and excluded them if they said yes. We then asked them to fill out a consent form. Then we introduced them to the ice cream topping adding task, and how to use an Xbox controller to teleoperate the arm. Participants had up to 3 minutes to practice the task before giving a demonstration. Each participant only had one chance to give a demonstration and could not retry. After giving a demonstration, the experimenter introduced participants to the replay evaluation task. The experimenter would introduce progress and scalar feedback in detail to reduce the difference in understanding of the signals among participants. During the replay, the arm would stop ten times. For each time, the experimenter would briefly explain progress and scalar feedback again, and ask participants for two signals in a random order.
V-C Non-Expert Demonstrations Are Multi-Policy and Noisy
@@
We collected 40 demonstrations from 40 participants. All trajectories are visualized in Figure 5. We also visualized the locations of the objects used in the experiment. The blue trajectories are successful demonstrations and the orange trajectories are failed demonstrations. We found that non-expert demonstrations are noisy and contain a variety of policies, even though 34 out of 40 are ultimately successful. This suggests that policies can be both successful and sub-optimal, which supports our intuition: assessing the quality of demonstrations by comparing is likely to lose information if there are many ”good enough” policies. We also observed that the noise in the demonstrations is not only teleoperating errors but also explorations. For instance, we observed that 14 participants slightly shook the topping jar to test if the gripper firmly held the jar when picking the jar, and 17 participants poured a few toppings out first to see if the jar was right above the ice cream when pouring the toppings. This highlights the importance of detecting unproductive behaviors, and suggests that noisy demonstrations from partially trained agents or perfect demonstrations with injected noise are inappropriate approximations of noisy human demonstrations: human demonstrations are “noisy” in a meaningful way. For the six failed demonstrations, three of them were because the topping jar was accidentally dropped while reaching the ice cream, and two of them were because the gripper did not successfully pick up the topping jar. The most common faulty cases are similar to the Failure scenario and the Unaware scenario we used in our online study, which confirmed the validity of the design of our online study.
@@ V-D Progress indicates task completions and is more consistent across participants than scalar feedback
We collected progress and scalar feedback from 34 participants (two participants’ progress and scalar feedback were excluded since all progress and scalar feedback they provided were 100). The results are shown in Figure 6. We plot the average progress and scalar feedback for all successful demonstrations, and progress and scalar individually for all failure demonstrations. We are not able to determine if progress from participants correctly describes the task completion rate since we do not have ground truth labels, but average progress for successful demonstrations did start from low and increased as the demonstrations were reproduced which is a strong indication that progress correlates with task completion rates.
We successfully identified all failed demonstrations by only looking at progress at the last step. We use 90 as the divide value, which is the average progress at the last step for successful demonstrations in our online study. If progress at the last step is less than 90, the task is incomplete, otherwise the task is completed. For all failed demonstrations, the end progress is less than 90 and the average progress at the end for successful demonstrations is 98.8. H3 is supported. Moreover, progress is significantly more consistent across participants even though their demonstrations are multiple policies and in different quality ($avg\_std_{progress}=11.6$, $avg\_std_{scalarfeedback}=25.5$, $p<0.001$, $BF>1000$). H4 is supported.
@@ V-E Progress allows the awareness of backtracking
Figure 7: 3D trajectory of one participant’s demonstration. The trajectories are colored with progress, and both scalar feedback and progress signals are indicated in a timeline above the figure. The brown circle represents the start point. The green star represents the endpoint. The blue square is the position of the topping jar. The participant reset the arm to the start position after demonstrating a few steps to adjust the grip point.
As mentioned in subsection V-C, participants would do explorations while demonstrating difficult parts of the task. We also observed that some participants went back a few steps to adjust the grip pose or the pouring position. For instance, one participant failed to pour the toppings into the ice cream because the grasp point was not optimal. Then the participant decided to reset the arm to the initial position, demonstrated the entire task from the beginning again, and succeeded. We plot the trajectory for that demonstration along with progress and scalar feedback, shown in Figure 7. As indicated by progress, the arm was reset to the initial position between step 4 and step 5, and the task was successfully completed afterward.
@@ VI Discussion
In this work, we investigated and closely defined an under-explored teaching signal, progress, and conducted three different studies to show the usefulness of progress across a variety of scenarios. We showed that giving progress is not hard and progress carries information beyond scalar feedback and other teaching signals in prior work. We expect that progress will be effective in many applications other than just using along with demonstrations.
Prevent Reward Hacking Progress could be a powerful signal to indicate reward hacking. Reward hacking is a phenomenon in which a learning agent learns to achieve high rewards by performing unintended actions instead of finishing the task. For instance, a cleaning robot gets a +1 reward every time it cleans a room. The robot, instead of cleaning one room and going to the next room, repetitively ejects dirt in a room, then cleans it, and thus achieves high rewards. Using progress, we can easily identify that the robot is not advancing towards task completion.
Inverse Reinforcement Learning We observed a similarity between progress functions and reward functions and expect that a progress function might be a representation of a reward function. We trained a reward function using demonstrations we collected in the public space study using Adversarial Inverse Reinforcement Learning [44], and calculated the rewards for each demonstration. We then calculated the Pearson correlation between the rewards of the demonstration and the progress, as well as the correlation between the rewards and the scalar feedback from each participant in Figure 8. We found that progress is strongly correlated with the learned reward function ($avg_{r}=0.83$) and the average correlation is significantly higher ($p<0.001$, $BF>1000$ ) than scalar feedback ($avg_{r}=0.19$).
Data Filtering and Ranking Progress could also be used to rank demonstrations. For example, a demonstration with a progress of 60 at the end should be ranked lower than a demonstration with a progress of 90. Moreover, progress can be applied as a data filter especially when the demonstrations are sub-optimal.
A key area for future work is to compare the model performance between the model trained using progress and the model trained using other types of human feedback. Training a reliable model from limited non-expert demonstrations and non-expert annotations is challenging but future work could expand our data with more demonstrations along with more types of human teaching signals such as preference.
@@ VII Conclusion
In conclusion, we defined progress in detail and found that progress could be used to describe completion degrees of a task, indicate if a task is complete, and be more consistent across users, without requiring extra workload or time compared to giving scalar feedback. We collected 40 non-expert demonstrations along with progress and scalar feedback, and released them as a dataset. We found that non-expert demonstrations are multi-policy and mostly successful, while noisy in a meaningful way. Our work suggests that progress is information-rich and is worth more attention to develop new methods to effectively leverage the novel information from progress.
ACKNOWLEDGMENT
The work described here was supported in part by the US National Science Foundation (IIS-2132887).